# Squish Jamming

### v1

@@ V Discussion
We established that large shape deformation may radically modify the mechanisms by which an assembly yields. At high density, far into the jammed state, the number of contacts becomes irrelevant, and their extent, coupled with frictional interactions, becomes crucial. The contact size is intrinsically linked to the geometrical deformations of the elements, so the shape also has a crucial role in the rigidity transition. Geometry and friction concur to generate an attractive force between rings under shear, and sufficient shear must be applied to overcome this effective adhesion. In the high friction limit, the shear must be enough to separate the rings completely to overcome the adhesion; hence, a simple geometric formula (Eq. (1)) sets the transition between repulsive and attractive interactions. Below this transition (when the interactions are attractive), structural rearrangements are strongly hindered, leading to a jammed assembly.
@@
However, these results are valid only within significant friction and shear limits. Our geometrical model implies that a frictionless assembly yields at an infinitesimal shear as $\gamma_{r}=\mu$. By considering only the interactions between two rings, we neglect collective effects, such as the caging [51], that would require a finite shear to yield even without friction. How caging is affected by significant shape changes is still an open question.
@@
Another striking property of the ring assembly is the drastic structure change between the fluid and solid state. Counterintuitively, the fluid phase features highly crystallised portions, while the solid remains in an amorphous structure. Similarly to the link between active matter and sheared granular media [52], one may also draw a parallel between this fluid-like crystal under shear and active crystals [53, 54].
@@
From an engineering perspective, assemblies of highly deformable elements combine frictional interactions and structural rearrangement to dissipate energy while keeping a significant elastic component [55]. Ring assemblies then have great potential as a tunable shock absorber, with the density serving as a knob to tune dissipation and response to a shock, keeping even a high dissipative potential in the solid and reversible state. To develop this axis, it would be essential to explore the 3D version of such assemblies and their response to inertial forces.
Acknowledgements.
Acknowledgments.
This work benefited from numerous enriching discussions with, but not limited to, A. Ikeda, N. Saito, S. Ishihara, Y. Lou, H. Ikeda, F. Lechenault, C. Scalliet, M. Henot, I. Cantat, B. Chakraborty, and B. Andreotti. We thank R. Tosaka for the construction of a prototype of the shear device. This work is supported in part by KAKENHI from the Japan Society for the Promotion of Science (Grant Nos.JP22KF0084, JP20H00128, JP19H05800, JP24K00593). S.P. acknowledges financial support from the Japanese Society for the Promotion of Science as a JSPS International Research Fellow.
Authors contribution.
S.P. and K.A.T. designed the experiments. S.P. performed the experiments, the numerical simulations, the data analysis, and conceived the model. All authors analysed and discussed the results and contributed to writing the paper.
@@ Acknowledgments.
The Zenodo repository 10.5281/zenodo.10961344, contains the experimental data for the force and from image analysis for ring assemblies and single ring experiments. The raw images of the first five cycles for $\gamma_{0}=0.40$ and $\Phi=1.00$ and the MATLAB code to process them. It also contains the FEM data and associated COMSOL files.
Figure 1: Loss and storage moduli across the parameter space. The first harmonic of the loss and storage moduli are shown in a and b, respectively. The moduli are normalised by the Young modulus of the rings, $E=1\,\mathrm{MPa}$.
Figure 2: Schematic for the geometrical predictions. a, First case to estimate the lateral displacement $x_{r}$ which for which the lateral force $f_{x}$ becomes repulsive ($f_{x}>0$). The contact is reduced to a single point with an orientation $\alpha$ given by the line linking the two intersection points of the half-circles. Assuming frictional Coulomb interactions, the amplitude of tangential $\mathbf{T_{t}}$ and normal $\mathbf{T_{n}}$ contact forces are linked by the friction coefficient $\mu$. The lateral force $f_{x}$ is the resultant of the contact forces projected on the lateral axis: $f_{x}=(\mathbf{T_{n}}+\mathbf{T_{t}})\cdot\mathbf{e_{x}}$. The expression for $x_{r}$ is then found by solving the equation $f_{x}(x=x_{r})=0$. This prediction is valid if the two rings are still in contact at $x=x_{r}=\mu(2r_{r}-\delta_{y})$. b, If the rings are not in contact at $x=x_{r}=\mu(2r_{r}-\delta_{y})$, we have to consider the second case where $f_{x}$ remains attractive until complete separation of the rings. In that case, $x_{r}$ is the no-overlap distance given by the initial contact extent $\ell_{c}$ after compression. c, To link the global assembly density $\Phi$ and the local compression $\delta_{y}$, we assume an isotropic and homogeneous compression of an assembly initially in a hexagonal packing. The density is estimated by considering a central ring and the hexagon with corners at the centre of its $6$ neighbours and by taking the ratio between the area of the ring inside the hexagon ($1$ full ring + $6$ third of rings) and the area of the hexagon. In the hexagonal lattice, the hexagon edge is $2r_{h}$ and reduces to $2r_{h}-\delta_{y}$ under compression, while the area of the rings is constant ($\Phi$ is estimated from the area of undeformed rings). d, Shear is a scaleless quantity defined as the ratio between the lateral displacement and the height of the system. The shear defined for the assembly as $\gamma_{=}X/T$ is then identical to the one in the local interaction experiments, $\gamma=x/(2r_{h}-\delta_{y})$.
Figure 3: The average coordination number is irrelevant for highly compressed assemblies. a, Average of the coordination number $Z_{r}$ of the assemblies in the steady state. The coordination number of a ring is defined as the number of touching neighbours. No clear tendency with $\gamma_{0}$, $\Phi$, or the fluid/solid regions is noticeable. b, Reproduction of Fig. 4e, showing that the relative value between the shear distance $x_{0}$ and the averaged maximal contact length $\langle{\ell_{c}}^{\mathrm{max}}\rangle$ marks a clear separation between fluid and solid-like regions. c, local average lateral displacement $x_{0}$ imposed to the assembly over the parameter space $(\gamma_{0},\Phi)$. It is a deterministic quantity computed following $x_{0}=\gamma_{0}2r_{h}\sqrt{\Phi_{h}/\Phi}$. d, For each tracked ring in the steady state, the contact length $\ell_{c}$ with each of its touching neighbours is measured (see Methods). At each time-step, the highest $\ell_{c}$ of each ring is averaged over all rings, and then all time-steps to give $\langle{\ell_{c}}^{\mathrm{max}}\rangle$. While $x_{0}$ mainly increases with the shear $\gamma_{0}$, $\langle{\ell_{c}}^{\mathrm{max}}\rangle$ is increasing with density $\Phi$. The two values coincide around the fluid-solid separation line predicted by $\gamma_{r}(\Phi)=\sqrt{\Phi/\Phi_{h}-1}$.

### v2

@@ Discussion
We established that large shape deformation may radically modify the mechanisms by which an assembly yields. At high density, far into the solid state, the number of contacts becomes irrelevant, and their extent, coupled with frictional interactions, becomes crucial. The extra internal degrees of freedom provided by shape changes help the rings to resist shear without rearranging by accommodating their shape. This is in sharp contrast with the classic picture of the jamming transition, where additional degrees of freedom would favor fluidization of the assembly, unless as many constraints are added to the system. We argue that this contrast originates from geometry and friction concurring to generate an attractive force between rings under shear, where sufficient shear must be applied to overcome this effective adhesion. In the high friction limit, the shear must be enough to separate the rings completely to overcome the adhesion; hence, a simple geometric formula (Eq. [1]) sets the transition between repulsive and attractive interactions. Below this transition (when the interactions are attractive), structural rearrangements are strongly hindered, then shape deformations take the upper hand and lead to a rigid assembly.
@@
However, these results are valid only within significant friction and shear limits. Our geometrical model implies that a frictionless assembly yields at an infinitesimal shear as $\gamma_{r}=\mu$. By considering only the interactions between two rings, we neglect collective effects, such as caging (32), that would require a finite shear to yield even without friction. How caging is affected by significant shape changes is still an open question. Investigating the low and intermediary friction regime is then an important challenge left for future studies. This may be tackled if one can experimentally and systematically vary the friction coefficient between two deformable interfaces, for instance, by modifying the external texture of the rings (57).
In addition, the nature of the observed rigidity transition is not fully elucidated. Further analysis of the particle dynamics via their mean-squared displacement in the steady state (SI Appendix, Fig. LABEL:S-fig_sup_transition) hints that the transition shares traits of both a jamming and a glass transition, with the shear amplitude playing the role of an effective temperature (58). However, the limited number of accessible cycles in the steady state hinders a conclusive characterization.
@@
Another striking property of the ring assembly is the drastic structure change between the yielded and solid states. Counter-intuitively, the yielded phase features highly crystallized portions, while the solid remains in an amorphous structure. Similarly to the link between active matter and sheared granular media (59), one may also draw a parallel between this yielding crystal under shear and active crystals (60, 61).
@@
From an engineering perspective, assemblies of highly deformable elements combine frictional interactions and structural rearrangements to dissipate energy while keeping a significant elastic component (62). Ring assemblies are then solid foams analogous with viscoelastic properties tunable by an external pressure field, with promising potential as a tunable shock absorber. A change in density around the rigidity transition point will greatly affect the material response to a shock. To develop this axis, it would be essential to explore the 3D version of such assemblies and their response to inertial forces, in particular, how granular inertial flows (63) are affected by shape changes.
[…]
Ring fabrication
Rings are fabricated from silicone elastomer (Elite Double 22, Zhermack) in a urethane mold (Clear Flex 50, Smooth-On). The ring’s material is chosen to remain elastic without permanent deformation, even under high and prolonged deformations. The mold material has been selected for its deformability, which eases demolding without tearing the rings, and its very low affinity with silicone, ensuring no adhesion with the rings and reusability. The rings have a thickness of $t=1.5\,\mathrm{mm}$, a midline radius $r=3.3\,\mathrm{mm}$ and height $h=10\,\mathrm{mm}$. To help with image analysis, a thin layer ($1\,\mathrm{mm}$) of dyed (Silc Pig Electric Orange, Smooth-On) silicone is applied on the inner portion of the ring surface (from $r-t/2$ to $r$). This extra layer does not affect the mechanics of the ring (the additional layer is not reproduced in FEM but still compares well with the experiments, Fig. 1B,C). The rings are slightly coated with talc powder to avoid adhesion and ensure reproducible frictional properties. See SI Appendix, Fig. LABEL:S-fig_sup_ring for an illustration of the protocol.
Shear setup
The shear cell comprises a fixed bottom arm and a movable top arm with movements constrained parallel to the bottom. The two sides of the cell freely rotate respectively to the top and bottom arms via joints of radius $40\,\mathrm{mm}$, and freely slide through the top arm via roller bearings to adjust their length under large oscillatory shear. The distance between the two side arms is $W=300\,\mathrm{mm}$, and the one between the top and bottom arm is $T=75\,\mathrm{mm}$. Both bottom and top arms are decorated by rigid (3D-printed in PLA and then glued) half disks of radius $r_{h}=r+t/2=4.05\,\mathrm{mm}$ (external radius of the rings). The whole setup rests on a fixed acrylic plate, and the ring assembly is confined vertically by a $3\,\mathrm{mm}$ thick acrylic plate blocked in the vertical direction but free to move laterally. The top arm imposes shear via a driving screw controlled by a stepper motor (AZM 46AC-TS10R, Oriental Motor). The motor imposes a cyclic horizontal motion $X=(X_{0}/2)\sin(2\dot{X_{0}}t/X_{0})$, with $X_{0}$ the displacement amplitude and $\dot{X_{0}}=1\,\mathrm{mm/s}$ the maximum shear speed, kept constant for all the amplitudes considered. The resulting oscillatory shear is $\gamma=X/T$. The density of the rings is calculated by $\Phi=N\pi{r_{h}}^{2}/A$, where $A=16500\,\mathrm{mm^{2}}$ is the area of the shear cell, evaluated by taking into account the size of the joints and the top and bottom granular boundary conditions. With the highest shear considered, this area varies between $16500\,\mathrm{mm^{2}}$ and $17500\,\mathrm{mm^{2}}$ within a shear cycle ($\sim 6\%$ of variations). Consequently, densities are defined with uncertainties of ±$\sim\pm 0.02$. The motor, force sensor, and camera are controlled and synchronized via a custom Labview program.
Force measurements
All forces reported here are measured with a flat six-axis (forces and moments along the three directions) force sensor (USX10-H10-500N-C, TecGihan) and transferred to a computer via 2 data acquisition cards (USB-6001, National Instrument). To measure the shear response of the assembly (Fig. 2B and $D$), the sensor is inserted between the top arm and the driving screw. The top arm is slightly lifted from the bottom acrylic plate to avoid measuring the frictional response of the setup. All six axes of the sensor are recorded at a rate of $20\,\mathrm{Hz}$. The first harmonic of the loss and storage moduli are computed by $G^{\prime}=(2/\pi\gamma_{0})\int_{0}^{2\pi}F_{y}(\theta)\sin\theta\mathrm{d}\theta/Wh$ and $G^{\prime\prime}=(2/\pi\gamma_{0})\int_{0}^{2\pi}F_{y}(\theta)\cos\theta\mathrm{d}\theta/Wh$ with $\theta=(t-t_{c})2\dot{\gamma_{0}}/\gamma_{0}$ the normalized time since the start of cycle $c$ (36). To measure the compression response of the rings (Fig. 1B,C) and their interactions (Fig. 5B,C), a distinct setup is used where positions are imposed manually via a micrometric stage and $100\,\mathrm{mm}$ long arm levers, and forces are measured via the moments output of the sensor. For the compression response, one free ring is confined between two rigid walls, while for the interaction experiments, a PLA 3D-printed part holding half of the rings into a groove clamps the pair of rings. For each interaction experiment, the six axes of the sensor are recorded continuously at $100\,\mathrm{Hz}$ while compression and shear are imposed via manually moving the micrometric stage. The displacement is made by step of $50\,\mathrm{\mu m}$ every $10\,\mathrm{s}$. For every step, the position is changed between 00 and $5\,\mathrm{s}$, and the reported values for this new position are averaged over the time interval between $7$ and $9\,\mathrm{s}$. The arm lever’s high but finite stiffness is measured in compression and shear and used to correct the displacement imposed on the rings.
Finite Element Method (FEM)
The mechanical tests using full 3D FEM were performed with the software COMSOL Multiphysics 6.1. We use symmetries to model only the necessary portions of the ring for the compression (Fig. 1B,C) and shear (Fig. 5) simulations. The rings are cylinders of height $h$ along the $z$ axis and of external radius $r_{h}$ in the $xy$ plane. Rings are compressed and sheared in the $xy$ plane, and we assume stress-free surfaces at $z=0$ and $z=h$. From symmetries, we model only half of the rings, from $z=0$ to $z=h/2$, with a free surface at $z=0$ and a mirror symmetry (no outward displacement) at $z=h/2$. The tetrahedral mesh size is fixed and set to a minimum of $\sqrt{rt}/5$, chosen to optimize convergence. The material is the Neo-Hookean model with Young modulus $E=1\,\mathrm{MPa}$ and Poisson’s ratio $\nu=0.45$. The Young modulus is determined by the best fit to the experimental data, while the Poisson ratio is taken just below $0.5$ as this silicone is nearly incompressible, but taking the limit value of $0.5$ brought non-convergence issues in our FEM simulations. The mechanical tests are performed via a stationary approach and by imposing displacements. For the compression tests (Fig. 1B,C), the ring is compressed between two flat and infinitely rigid blocks, and contacts are solved via the Nitsche method. For the interaction tests (Fig. 5), two facing rings are cut again along the $x$ direction to form semi-circles. To clamp the rings, no displacement in the $xz$ plane is imposed on the cut surfaces of the top ring. The position of the bottom ring, first along $y$ for compression and then $x$ for shearing, is also controlled by imposing the displacement of its cut surfaces. The contacts are also solved using the Nitsche method, but this time adding a Coulomb frictional interaction term with a unique friction coefficient $\mu$. The resulting shear force $f_{x}$ is then obtained by integrating the $\sigma_{yx}$ component of the stress on the surfaces at $y=0$ of the bottom ring, where the shear displacement is imposed.
Image acquisition and analysis
Pictures are acquired via a CMOS camera (EMVC-CB1400C3 with an EMVL-MP814 lens, Misumi) placed on top of the assembly. Indirect lighting is obtained by shining a light-emitting diode panel on a whiteboard. Irrespective of shear amplitude, $15$ pictures per shear cycle are taken ($30$ pictures for $(\gamma_{0},\Phi)=(0.400,0.94),(0.533,0.94),(0.533,1.00)$ where tracking is more challenging), and a $5\,\mathrm{s}$ pause between the cycles is implemented to ensure that each cycle’s first and last pictures correspond to $\gamma=0$. The pictures consist of $4608\times 1300\,\mathrm{pixels}$, resulting in a resolution of $13.05\,\mathrm{pixel/mm}$. Then, a custom-made Matlab script processes all the pictures. Using the orange-colored inner part of the rings, they are detected and separated via color and saturation thresholding. A group of pixels within its contour is attributed to each detected ring and, for instance, used to represent ring-specific quantities via color coding, as in Fig. 2A. The geometrical center of the pixels defines then the position of the ring, while its shape, or contour, is measured via a circle Fourier decomposition (29). The polar coordinates $R(\theta)$ of the outermost pixels of the contour are measured with the geometrical center of the ring as the origin. Using the fast Fourier transform, the Fourier coefficients $(a_{k},b_{k})$ of the contour are measured such that $R(\theta)=r_{h}\sum_{k=0}^{M}(a_{k}\cos\theta+b_{k}\sin\theta)$ up to $M=5$. In addition to the position and shape of the ring, we also detect their neighbors via two methods. First, we find topological neighbors via their position by performing the Delaunay tessellation. Based on the resulting Voronoi diagram, we defined Voronoi neighbors as rings sharing a common edge and no farther than $2.5r$. From this Voronoi neighborhood, quantities such as the orientational bond order parameter $\Psi_{6}$ are computed (see Fig. 3). The other neighborhood detection method aims to detect the rings in contact. A given ring’s contour is dilated by $8\,\mathrm{pixels}$, and its touching neighbors are those whose contour overlaps. For these touching rings, we then compute the bond vectors from their geometrical center and the contact extents for Fig. 5F. To measure the extent of the contacts $\ell_{c}$, we compute the long axis of the overlapping contours. Indeed, each overlap has an approximate ellipsoidal shape with its long axis along the contacting surfaces (see ring overlaps in SI Appendix, Fig. LABEL:S-fig_sup_schemB). This measure overestimates the actual contact length because of the dilatation of the central ring, and we found that multiplying by $0.7$ gives a visually convincing measure. Apart from $\Psi_{6}$, while we focused on the average (over all rings in the steady state) of the quantities of interest, their distribution is also available in SI Appendix, Fig. LABEL:S-fig_sup_distri.
Geometrical and topological strains
To distinguish quantitatively the geometrical deformations (change of shape) from the structural or topological ones (change in neighbors), we follow the method described in great detail in the Appendix section of Ref. (16). Initially developed for epithelial tissue development, this strain quantification can be simplified and applied to our ring assembly by taking into account the fact that ring assemblies are not confluent, and that the topological deformations arise only from neighbors’ changes, in contrast to living cells which may also divide. The geometric and topological strains, denoted respectively by $S^{*}$ and $T^{*}$, are computed between two consecutive time frames from the variations (length, orientation, existence) of the bonds between rings. The computation is based on the $2\times 2$ texture tensors of the assembly, defined by $\mathbf{M}=\frac{1}{2}\sum_{B_{\text{tot}}}\vec{L}\otimes\vec{L}$ and $\mathbf{m}=\frac{1}{2}\sum_{b_{\text{tot}}}\vec{l}\otimes\vec{l}$, for the earlier and later consecutive frames, respectively. Here, $B_{\text{tot}}$ (resp. $b_{\text{tot}}$) is the set of all bond vectors $\vec{L}$ (resp. $\vec{l}$) in the assembly in the earlier (resp. later) frame, and $\vec{L}\otimes\vec{L}=\vec{L}\vec{L}^{T}$ is the tensor product of $\vec{L}$ by itself. Then we compute the deformation gradient tensor $\mathbf{F}=(\mathbf{F_{0}}\mathbf{F_{1}})^{1/2}$, with $\mathbf{F_{0}}=(\sum_{b_{\text{cons}}}\vec{l}\otimes\vec{L})(\sum_{b_{\text{cons}}}\vec{L}\otimes\vec{L})^{-1}$ and $\mathbf{F_{1}}=(\sum_{b_{\text{cons}}}\vec{l}\otimes\vec{l})(\sum_{b_{\text{cons}}}\vec{L}\otimes\vec{l})^{-1}$, where $b_{\text{cons}}$ is the set of the conserved bonds, i.e., the bonds that exist in both frames. We also compute the strain correction tensor, $\mathbf{\Psi}=\mathbf{F}(\mathbf{M_{c}}\mathbf{F}^{T}-\mathbf{F}^{T}\mathbf{M_{c}})$, with $\mathbf{M_{c}}=\frac{1}{2}\sum_{b_{\text{cons}}}\vec{L}\otimes\vec{L}$. For the contribution from the non-conserved bonds, we evaluate the topological tensor defined by $\mathbf{T}=\frac{1}{2}\sum_{b_{\text{topo}}}\vec{l}\otimes\vec{l}-\frac{1}{2}\sum_{B_{\text{topo}}}\vec{L}\otimes\vec{L}$, where $b_{\text{topo}}$ (resp. $B_{\text{topo}}$) is the set of the bonds that appear (resp. disappear) between the two frames (only due to rearrangements in our case, as opposed to living cells). Finally, we nondimensionalize each tensor $\mathbf{Q}$ by $\mathbf{\tilde{Q}}=\frac{1}{4}(\mathbf{Q}\mathbf{M_{c}}^{-1}+\mathbf{M_{c}}^{-1}\mathbf{Q})$ and obtain the desired strains by $\mathbf{S^{*}}=\frac{|B_{\text{tot}}|}{|b_{\text{tot}}|}\mathbf{\tilde{m}}-\mathbf{\tilde{M}}-\mathbf{\tilde{\Psi}}$ and $\mathbf{T^{*}}=\frac{|b_{\text{tot}}|-|B_{\text{tot}}|}{|b_{\text{tot}}|}\mathbf{\tilde{m}}-\mathbf{\tilde{T}}$, where $|B_{\text{tot}}|$ (resp. |$b_{\text{tot}}$|) is the number of the elements of the corresponding set, i.e., the total number of the bonds in the earlier (resp. later) frame. Note that the total strain can be simply expressed, approximately, as $\mathbf{S^{*}}+\mathbf{T^{*}}\simeq\frac{1}{2}(\mathbf{F}\mathbf{F}^{T}-\mathbb{1})$ (valid for $\norm{\mathbf{S}^{*}+\mathbf{T}^{*}}\ll 1$). To sum up, by measuring the bonds in both frames and identifying the conserved, created and eliminated ones, we can compute the geometrical and topological strain tensors, $\mathbf{S^{*}}$ and $\mathbf{T^{*}}$, respectively. To obtain simpler quantifiers of the geometrical and topological contributions, here we take the norm (the square root of the sum of the squared elements of the matrix) of these strain tensors and sum them over the $6$ shear cycles in the steady state. The resulting quantities, $S_{c}=\int_{t}\norm{\mathbf{S^{*}}}\mathrm{d}t$ and $T_{c}=\int_{t}\norm{\mathbf{T^{*}}}\mathrm{d}t$, are the scalar quantities shown in Fig. 4E,F.
@@
Supplementary videos, experimental raw force data and data processed by image analysis of the ring assemblies, for all the parameters explored; FEM data and processed experimental data for the single ring compression and interactions; COMSOL files to run the FEM simulations and the associated Neo-Hookean material law; and raw images of the five first cycles for $\gamma_{0}=0.40$ and $\Phi=1.00$, with a MATLAB code to process the images and extract the relevant quantities (each category comes with a dedicated README.txt that details its content) have been deposited in Zenodo (64). Some study data are available: due to size restrictions, all the experimental raw images of the ring assemblies are not included in the repository. However, the entire image dataset can be shared with interested researchers upon request.
[…]
This work benefited from numerous enriching discussions with, but not limited to, A. Ikeda, O. Dauchot, N. Saito, S. Ishihara, Y. Lou, H. Ikeda, F. Lechenault, C. Scalliet, M. Henot, I. Cantat, B. Chakraborty, and B. Andreotti. We thank R. Tosaka for the construction of a prototype of the shear device. This work is supported in part by KAKENHI from the Japan Society for the Promotion of Science (Grant Nos. JP22KF0084, JP20H00128, JP19H05800, and JP24K00593). S.P. acknowledges financial support as a Japan Society for the Promotion of Science International Research Fellow.
[…]